Circular No. 8100

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1550–564 P. Dubath, INTEGRAL Science Data Center and Geneva Observa-

tory; M. Revnivtsev, Space Research Institute, Moscow, and Max-Planck-Institut für Astrophysik, Garching; P. Goldoni, Centre d’Etude Atomique and Service d’Astrophysique, Saclay; A. von Kienlin, Max-Planck-Institut für Extraterrestrische Physik; N. Lund, Danish Space Research Institute, Copenhagen; S. Grebenev, Space Research Institute, Moscow; and E. Kuulkers, Research and Science Support Department, ESTEC, Noordwijk, report a new x-ray outburst of the transient source XTE J1550–564 (cf. IAUC 7010), detected with the IBIS and SPI instruments onboard the INTEGRAL satellite. Between Mar. 24.2 and 25.1 UT, the average fluxes were ∼ 150 mCrab in the energy band 15–40 keV and ∼ 250 mCrab in the band 40–100 keV. The photon index of the spectrum is ∼ 1.5 in the energy band 20–150 keV. Observations at other wavelengths are encouraged.

η CARINAE

R. S. Levenhagen, A. Damineli, N. V. Leister, and Z. Abraham, Universidade de São Paulo and the State of São Paulo Research Foundation, report on spectroscopy of η Car made on Mar. 15.2 UT at Laboratório Nacional de Astrofisica. The ‘red’ peak in the He I line at 1083 nm, which was stronger than the ‘violet’ peak during the last five years, is now fainter and decreasing rapidly. This is a clear sign that the ‘spectroscopic event’ predicted for late June 2003 is under way (Damineli et al. 2000, Ap.J. 528, L101); based on this single observation, a lower limit of 2005 days is derived for the periodicity of such events.

COMETS C/2002 W1, C/2002 W2, C/2002 W3, C/2002 W4 (SOHO)

Further to IAUC 8099, D. Hammer reports measurements for additional Kreutz sungrazing comets found on SOHO C2 website images by X.-M. Zhou (C/2002 W1), R. Kracht (C/2002 W2, C/2002 W3), and M. Meyer (C/2002 W4). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 W1 Nov. 16.642  15^{h}27.^{m}6  −20^{o}36′  2003-F16
C/2002 W2  17.775  15 30.7  −20 43  2003-F16
C/2002 W3  18.742  15 33.8  −20 52  2003-F18
C/2002 W4  19.535  15 37.0  −21 02  2003-F18
2003 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green